7. A skier with a mass of 62 kg is sliding down a snowy slope. Find the coefficient of kinetic friction for the skier if friction is known to be 45.0 N


1
Expert's answer
2021-09-23T08:33:10-0400

The force friction is


"f=\mu mg\sin\theta,\\\space\\
\mu=\frac{f}{mg\sin\theta}=\frac{0.0741}{\sin\theta},"

where theta is the angle of the slope.